# HEIST: A Graph Foundation Model for Spatial Transcriptomics and Proteomics Data

### v2

@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has revolutionized our ability to study gene expression at the resolution of individual cells, offering data-driven insights into biology. The complexity of these datasets has fueled the development of machine learning methods for modeling cellular diversity, predicting cell states, and imputing or denoising gene expression values  (Cui et al., 2024; Hu et al., 2021; Grønbech et al., 2020; Wen et al., 2023a; Bravo Gonzalez-Blas et al., 2024; Dijk et al., 2017). However, a limitation of scRNA-seq is the lack of spatial context of cells within tissues, which is important for understanding processes such as tissue organization, microenvironment interactions and how gene co-expression influences tissue-level behaviors. Spatial transcriptomics is an emerging technology that bridges this gap by preserving the physical locations of gene expression measurements, enabling a holistic study of tissue architecture, cell-cell communication, and their aberrations in the tumor context (Xiao & Yu, 2021; Rodriques et al., 2019). Similarly, spatial proteomics assays were developed to directly capture protein abundance, expression and signaling pathways, offering a complementary layer of biological insight. Despite advances in spatial omics, datasets remain limited by low throughput, platform and tissue heterogeneity, and scarce labels, often necessitating dataset-specific models (e.g., MIBI (Angelo et al., 2014), Imaging CyTOF). A foundation model trained on diverse spatial omics can address these challenges by learning generalized representations across tissues, organs, settings, and technologies, enabling strong performance on downstream tasks even with limited data.
@@
Prior foundation models such as scGPT (Cui et al., 2024), scFoundation (Hao et al., 2023), and CellPLM (Wen et al., 2023a) either neglect cell–cell structure or are limited to predefined gene sets, hindering generalization to unseen genes and proteins. Graph-based methods like GraphST (Long et al., 2023) and STAGATE (Dong & Zhang, 2022) (Explained in detail in Appendix A) capture spatial neighborhoods but remain task-specific and non-transferable. While transformer-based models like scGPT-spatial (Wang et al., 2025) treat all genes as fully connected graph and fail to model the inductive bias of co-expression. HEIST overcomes these gaps by associating each cell with a co-expression network that interacts with its spatial neighbors, incorporating microenvironmental cues into representations.
@@ 2 Method
Hierarchical Graph Construction. As shown in Figure A1, we first preprocess the data by removing outliers, normalizing gene expression, and retaining highly variable genes. Then we apply MAGIC (Dijk et al., 2017) to denoise gene expression values and reduce dropout noise. To build the gene co-expression networks, we first subset the cells based on cell-types using provided annotations or Leiden clustering. Following this, we compute pairwise mutual information between denoised genes within each type, and connect gene pairs above a threshold $\tau$. This results in total of $|{\mathcal{T}}|$ gene co-expression networks with mutual information prior. We create the spatial cell–cell graph by computing Voronoi polygons from cell coordinates and connecting cells in adjacent polygons. We then connect each cell with the gene co-expression network of that cell-type. The resulting outputs are a spatial graph ${\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}})$ and a set of gene co-expression networks ${{\mathcal{G}}_{g}^{t_{k}}({\mathcal{V}},{\mathcal{E}}_{t_{k}},{\mathbf{X}}_{k})}_{k=1}^{|{\mathcal{C}}|}$. We discuss the graph creation at length in Appendix B. Given these graphs, HEIST computes cell embeddings and gene embeddings ${\mathbf{Z}}_{c}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ and ${\mathbf{Z}}_{g}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$, such that
@@
Figure 2: HEIST Architecture
@@
where $\tilde{{\mathbf{H}}}_{c}^{(l)^{\text{repeat}}}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$ represents the cell embeddings repeated $|{\mathcal{V}}|$ times to align with the gene embeddings in each cell. Each gene receives information from its parent cell, enabling spatial context to modulate gene-level representations. Conversely, $\bar{{\mathbf{H}}}_{g}^{(l)}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ is obtained by aggregating the gene embeddings within each cell: $\bar{{\mathbf{H}}}_{g}^{(l)}=[\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{1}),\ldots,\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{|{\mathcal{C}}|})]$, where $\texttt{AGG}(\cdot)$ can be aggregation function such as MEAN pooling or differential pooling (Ying et al., 2018). This allows the cell embedding to be informed by the internal transcriptional state of the cell.
@@
where, HEIST-Decoder is a 3-layer GIN network (Xu et al., 2018).
@@
$\ell_{c\leftrightarrow g}$ $=\log\tfrac{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}}{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}+\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{k}_{g})/\tau}},$
$\mathcal{L}_{\text{contrastive}}$ $=\sum_{i,j\in{\mathcal{P}}}\!\big(\ell_{c\leftrightarrow c}({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})+\ell_{c\leftrightarrow g}({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})\big)+\sum_{(p,q)\in{\mathcal{V}}^{2}}\!\ell_{g\leftrightarrow g}({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q}).$
@@
Final loss. Contrastive learning structures the latent space to emphasize biological similarities and differences, promoting better separation of cell types and gene programs. In contrast, masked autoencoding ensures that embeddings retain rich information content necessary for reconstructing gene expression and spatial locations. Together, they prevent trivial or collapsed representations and produce embeddings that are both discriminative and information-rich. Hence, the final objective is a weighted sum of the contrastive and autoencoding losses, along with an orthogonality regularization which encourages the embedding dimensions to be decorrelated, promoting diverse and non-redundant representations (Zhang et al., 2021):
@@
where $\lambda$ is a regularization weight, and $\gamma$ is a learnable scalar that dynamically balances two terms.
@@ 3 Experiments
Pretraining Datasets. HEIST is trained on a large and diverse collection of high-resolution spatial transcriptomics datasets, primarily generated using single-cell technologies such as MERFISH and Xenium. The pretraining dataset comprises 22.3M cells from 124 tissue slices across 15 organs [cf. Figure 1(A)], including 13.3M cells from 10x Genomics  [^1] https://www.10xgenomics.com/, 8.7M from Vizgen  [^2] https://vizgen.com/, and 360K cells from the Seattle Alzheimer’s Brain Atlas  [^3] https://portal.brain-map.org/explore/seattle-alzheimers-disease. The large scale and diversity of the dataset improve the reliability and robustness of learned representations, enabling better transferability to downstream tasks across varying biological contexts and technologies. A detailed breakdown of datasets, descriptions, and sources is provided in Appendix D, and we provide a spreadsheet of the dataset details in the supplementary material.
Experimental setup. We provide pretraining hyperparameters in Appendix Table A1. HEIST was pretrained on 4 NVIDIA L40s GPUs (40GB each), with each epoch taking approximately 3 hours. Although the maximum number of epochs was set to 20, early stopping based on validation loss typically halted training around Epoch 5 or 6. For downstream evaluations, we assess HEIST in both zero-shot and fine-tuning settings. In the zero-shot setting, the pretrained model is directly evaluated on unseen spatial transcriptomics datasets to assess generalization without further training. For fine-tuning, we first extract embeddings from the frozen model and either train an MLP prediction head or fine-tune the decoder. We perform each experiment 5 times provide mean and standard deviation, except in Charville and UPMC datasets where we use the split the data using method from (Wu et al., 2022). The code is available at https://anonymous.4open.science/r/HEIST-E3F8.
@@
Cell clustering is critical for discovering novel cell types and understanding how microenvironmental factors shape cellular behavior, particularly in tumor microenvironments. An expressive model should not only cluster cells but also reveal microenvironment-driven subclusters, providing insights into spatially informed cell populations. Clustering is performed using frozen embeddings, and performance is evaluated on datasets two spatial transcriptomics datasets SEA-AD (Gabitto et al., 2024) and Merfish Lung Cancer (Chen et al., 2024), and three proteomics datasets Charville, UPMC, and DFCI (Wu et al., 2022), using normalized mutual information (NMI). Cell type annotation classifies cells into known biological categories, enabling functional interpretation of cellular diversity. HEIST embeddings are extracted from the frozen encoder, and an MLP classifier is trained on labeled datasets including SEA-AD, Charville, UPMC, DFCI, and MERFISH lung cancer, with performance evaluated using F1 score.
@@
Clinical outcome prediction aims to classify entire tissues, predicting outcomes such as immunotherapy response, treatment outcomes, remission status, and placenta condition. This task is essential for clinical decision-making and understanding disease progression. HEIST is evaluated on datasets including proteomics data Charville (Colon), UPMC (Neck), and DFCI (Neck) (Wu et al., 2022) collected using CODEX, skin cancer data (Ptacek et al., 2021) collected using MIBI, and spatial transcriptomics data of placenta collected using Xenium. Predictions are made using frozen cell embeddings and an MLP classifier, and we report AUC-ROC. Gene imputation recovers missing or noisy gene expression values, a common issue in spatial transcriptomics due to measurement limitations. We perform gene imputation by predicting masked gene values, using stratified sampling based on gene sparsity following the approach of Avşar & Pir (2023). This task is evaluated in both zero-shot and by fine-tuning the decoder, reporting Pearson correlation between predicted and true gene expression values. We explain these tasks in further details in Appendix D.2.
@@
Baselines. We benchmark HEIST against a diverse set of baselines, including graph-based spatial models, single-cell foundation models, and recent spatial foundation approaches. In particular, we include STAGATE (Dong & Zhang, 2022) and GraphST (Long et al., 2023), which capture local spatial relationships but struggle to generalize across datasets. We also evaluate against scFoundation (Hao et al., 2023), a large-scale single-cell foundation model that ignores spatial context. We also include comparisons with recent spatial foundation models CellPLM (Wen et al., 2023a), NoVAE (Blampey et al., 2024) and scGPT-spatial (Wang et al., 2025), which incorporate spatial information but do not explicitly model hierarchical gene-cell interactions. Finally, for the gene imputation task, we compared against MAGIC a denoising method based on graph diffusion. For proteomics datasets such as UPMC, baselines like scGPT-spatial and CellPLM were originally pretrained on fixed gene vocabularies, supporting only 6 of the 28 markers available. To enable comparison, we report two settings: (i) unaligned, where models process only the subset of overlapping genes, and (ii) aligned, where we manually remap marker identities to the closest supported genes following prior work. It is important to note that scGPT pretrains on single-cell spatial trasncritomics data mixed with Visium (spot-level), while CellPLM combines scRNA-seq with single-cell spatial transcriptomics. HEIST, in contrast, is trained exclusively on single-cell resolution spatial transcriptomics. However, there is major overlap in the single-cell spatial transcriptomics portion of training data between HEIST and the baselines.
@@ 3.1 Results
Table 2: Performance on clinical outcome prediction. We classify cancer outcome, cancer remission, treatment response and placental conditions. Results on the right of vertical are proteomics. NEM stands for "not enough markers" error from Novae when it can not match enough markers. *Results from (Wu et al., 2022)
@@
Ablation of components. Table 4 shows that removing any key component of HEIST leads to performance degradation across tasks, confirming their importance. Hierarchical modeling and spatial information are most critical, with their removal causing the largest drops. Pre-training is crucial, especially for clinical outcome prediction due to the skewed label distribution in these datasets. Cross-level message passing and contrastive learning significantly improve cell classification, while the masked autoencoder (MAE) is crucial for gene imputation and classification.
@@ Appendix A Related works
Hierarchical graph neural networks. Hierarchical graph neural networks operate on graphs with multiple levels, where each level represents a distinct granularity or type of information. MSMGN (Fortunato et al., 2022), HOOD (Grigorev et al., 2023), and BSMS-GNN (Cao et al., 2023) perform cross-level message passing across levels, but the levels are of same modality at different resolutions (e.g., coarsened versions of spatial graphs), rather than integrating graphs of distinct modalities and they are geared towards propagating signals globally. HIGH-PPI (Gao et al., 2023) introduces a hierarchical graph where one level encodes protein-protein interactions and the other encodes residue-level interactions. However, it does not include cross-level message passing, which limits the ability to combine information across biological levels. In contrast, HEIST works information from different modalities while integrating information across levels.
@@
Graph-based Spatial Transcriptomics methods. Graph-based spatial transcriptomics methods construct cell graphs based on spatial proximity and apply graph neural networks to model spatial relationships. SpaGCN (Hu et al., 2021), CCST (Li et al., 2022), STAGATE (Dong & Zhang, 2022), ConST (Zong et al., 2022), and GraphST (Long et al., 2023) train graph neural networks in un/self-supervised manner by using techniques such as deep graph infomax (Veličković et al., 2018), iterative clustering (Braun et al., 2022), and auto-encodings task on gene expression (Fang et al., 2024). Unlike HEIST, these models do not generalize to other tissues without retraining and do not incorporate hierarchical modeling across genes and cells.
@@
scRNA-seq and Spatial Transcriptomics Foundation models. Foundation models for single-cell and spatial transcriptomics aim to pretrain general-purpose representations that can be transferred across datasets and tasks. scFoundation (Hao et al., 2023), geneformer (Theodoris et al., 2023), and scGPT (Cui et al., 2024) extend transformer architectures to model scRNA-seq data by treating gene expression profiles as sequences. However, these models assume an ordering over genes and do not explicitly capture cell-cell relationships or spatial information. scGPT-spatial (Wang et al., 2025) adapts scGPT to spatial transcriptomics. CellPLM (Wen et al., 2023a) incorporates cell-cell interactions in pertaining and adds a gaussian mixture prior to overcome data limitations. But CellPLM and scGPT-spatial operate over a fixed set of genes and do not model hierarchy. While transformer based models like scGPT-spatial can be considered as graph neural networks over fully conneceted graphs Joshi (2025), they most biological networks are far from being fully connected, and explicitly modeling sparse, biologically plausible interactions is critical and enabled by graph-based learning. Existing foundation models thus either neglect spatial dependencies, gene co-expression structure, or both, limiting their ability to capture the context-specific nature of spatial transcriptomics. Furhtermore, their reliance in on gene embeddings hinders their generalization to proteomics data. In contrast, HEIST hierarchically models both this information and computes expressive and generalizable representations by working over gene co-expression networks.

### v3

@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has revolutionized our ability to study gene expression at the resolution of individual cells, offering data-driven insights into biology. The complexity of these datasets has fueled the development of machine learning methods for modeling cellular diversity, predicting cell states, and imputing or denoising gene expression values  (cui2024scgpt; spagcn; scvae; wen2023cellplm; bravo2024single; dijk2017magic). However, a limitation of scRNA-seq is the lack of spatial context of cells within tissues, which is important for understanding processes such as tissue organization, microenvironment interactions and how gene co-expression influences tissue-level behaviors. Spatial transcriptomics is an emerging technology that bridges this gap by preserving the physical locations of gene expression measurements, enabling a holistic study of tissue architecture, cell-cell communication, and their aberrations in the tumor context (xiao2021tumor; rodriques2019slide). Similarly, spatial proteomics assays were developed to directly capture protein abundance, expression and signaling pathways, offering a complementary layer of biological insight. Despite advances in spatial omics, datasets remain limited by low throughput, platform and tissue heterogeneity, and scarce labels, often necessitating dataset-specific models (e.g., MIBI (mibi), Imaging CyTOF). A foundation model trained on diverse spatial omics can address these challenges by learning generalized representations across tissues, organs, settings, and technologies, enabling strong performance on downstream tasks even with limited data.
@@
Prior foundation models such as scGPT (cui2024scgpt), scFoundation (hao2023large), and CellPLM (wen2023cellplm) either neglect cell–cell structure or are limited to predefined gene sets, hindering generalization to unseen genes and proteins. Graph-based methods like GraphST (long2023spatially) and STAGATE (dong2022deciphering) (Explained in detail in Appendix A) capture spatial neighborhoods but remain task-specific and non-transferable. While transformer-based models like scGPT-spatial (wang2025scgpt) treat all genes as fully connected graph and fail to model the inductive bias of co-expression. HEIST overcomes these gaps by associating each cell with a co-expression network that interacts with its spatial neighbors, incorporating microenvironmental cues into representations.
@@ 2 Method
Hierarchical Graph Construction. As shown in Figure A1, we first preprocess the data by removing outliers, normalizing gene expression, and retaining highly variable genes. Then we apply MAGIC (dijk2017magic) to denoise gene expression values and reduce dropout noise. To build the gene co-expression networks, we first subset the cells based on cell-types using provided annotations or Leiden clustering. Following this, we compute pairwise mutual information between denoised genes within each type, and connect gene pairs above a threshold $\tau$. This results in total of $|{\mathcal{T}}|$ gene co-expression networks with mutual information prior. We create the spatial cell–cell graph by computing Voronoi polygons from cell coordinates and connecting cells in adjacent polygons. We then connect each cell with the gene co-expression network of that cell-type. The resulting outputs are a spatial graph ${\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}})$ and a set of gene co-expression networks ${{\mathcal{G}}_{g}^{t_{k}}({\mathcal{V}},{\mathcal{E}}_{t_{k}},{\mathbf{X}}_{k})}_{k=1}^{|{\mathcal{C}}|}$. We discuss the graph creation at length in Appendix B. Given these graphs, HEIST computes cell embeddings and gene embeddings ${\mathbf{Z}}_{c}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ and ${\mathbf{Z}}_{g}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$, such that
@@
Figure 2: HEIST Architecture. HEIST Layer block presented in Figure A2.
@@
where $\tilde{{\mathbf{H}}}_{c}^{(l)^{\text{repeat}}}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$ represents the cell embeddings repeated $|{\mathcal{V}}|$ times to align with the gene embeddings in each cell. Each gene receives information from its parent cell, enabling spatial context to modulate gene-level representations. Conversely, $\bar{{\mathbf{H}}}_{g}^{(l)}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ is obtained by aggregating the gene embeddings within each cell: $\bar{{\mathbf{H}}}_{g}^{(l)}=[\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{1}),\ldots,\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{|{\mathcal{C}}|})]$, where $\texttt{AGG}(\cdot)$ can be aggregation function such as MEAN pooling or differential pooling (ying2018hierarchical). This allows the cell embedding to be informed by the internal transcriptional state of the cell.
@@
where, HEIST-Decoder is a 3-layer GIN network (xu2018powerful).
@@
$\ell_{c\leftrightarrow g}$ $=\log\tfrac{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}}{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}+\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{k}_{g})/\tau}},\quad\mathcal{L}_{\text{contrastive}}=\sum_{i,j\in{\mathcal{P}}}\!\big(\ell_{c\leftrightarrow c}+\ell_{c\leftrightarrow g}\big)+\sum_{(p,q)\in{\mathcal{V}}^{2}}\!\ell_{g\leftrightarrow g}.$
@@
Final loss. Contrastive learning structures the latent space to emphasize biological similarities and differences, promoting better separation of cell types and gene programs. In contrast, masked autoencoding ensures that embeddings retain rich information content necessary for reconstructing gene expression and spatial locations. Together, they prevent trivial or collapsed representations and produce embeddings that are both discriminative and information-rich. Hence, the final objective is a weighted sum of the contrastive and autoencoding losses, along with an orthogonality regularization which encourages the embedding dimensions to be decorrelated, promoting diverse and non-redundant representations (cca-ssg):
@@
where $\lambda$ is a regularization weight, $\sigma$ is sigmoid function to balance the loss terms, and $\gamma$ is a learnable scalar that dynamically balances two terms.
@@ 3 Experiments
Pretraining Datasets. HEIST is trained on a large and diverse collection of high-resolution spatial transcriptomics datasets, primarily generated using single-cell technologies such as MERFISH and Xenium. The pretraining dataset comprises 22.3M cells from 124 tissue slices across 15 organs [cf. Figure 1(A)], including 13.3M cells from 10x Genomics, 8.7M from Vizgen, and 360K cells from the Seattle Alzheimer’s Brain Atlas. The large scale and diversity of the dataset improve the reliability and robustness of learned representations, enabling better transferability to downstream tasks across varying biological contexts and technologies. A detailed breakdown of datasets, descriptions, and sources is provided in Appendix D, and we provide a spreadsheet of the dataset details in the supplementary material.
Experimental setup. We provide pretraining hyperparameters in Appendix Table A1. HEIST was pretrained on 4 NVIDIA L40s GPUs (40GB each), with each epoch taking approximately 3 hours. Although the maximum number of epochs was set to 20, early stopping based on validation loss typically halted training around Epoch 5 or 6. For downstream evaluations, we assess HEIST in both zero-shot and fine-tuning settings. In the zero-shot setting, the pretrained model is directly evaluated on unseen spatial transcriptomics datasets to assess generalization without further training. For fine-tuning, we first extract embeddings from the frozen model and either train an MLP prediction head or fine-tune the decoder. We perform each experiment 5 times provide mean and standard deviation, except in Charville and UPMC datasets where we use the split the data using method from (wu2022space). The code is available at https://anonymous.4open.science/r/HEIST-E3F8.
@@
Cell clustering is critical for discovering novel cell types and understanding how microenvironmental factors shape cellular behavior, particularly in tumor microenvironments. An expressive model should not only cluster cells but also reveal microenvironment-driven subclusters, providing insights into spatially informed cell populations. Clustering is performed using frozen embeddings, and performance is evaluated on datasets two spatial transcriptomics datasets SEA-AD (sea-ad) and Merfish Lung Cancer (chen_2024_11198494), and three proteomics datasets Charville, UPMC, and DFCI (wu2022space), using normalized mutual information (NMI). Cell type annotation classifies cells into known biological categories, enabling functional interpretation of cellular diversity. HEIST embeddings are extracted from the frozen encoder, and an MLP classifier is trained on labeled datasets including SEA-AD, Charville, UPMC, DFCI, and MERFISH lung cancer, with performance evaluated using F1 score.
@@
Clinical outcome prediction aims to classify entire tissues, predicting outcomes such as immunotherapy response, treatment outcomes, remission status, and placenta condition. This task is essential for clinical decision-making and understanding disease progression. HEIST is evaluated on datasets including proteomics data Charville (Colon), UPMC (Neck), and DFCI (Neck) (wu2022space) collected using CODEX, skin cancer data (melanoma-data) collected using MIBI, and spatial transcriptomics data of placenta collected using Xenium. Predictions are made using frozen cell embeddings and an MLP classifier, and we report AUC-ROC. Gene imputation recovers missing or noisy gene expression values, a common issue in spatial transcriptomics due to measurement limitations. We perform gene imputation by predicting masked gene values, using stratified sampling based on gene sparsity following the approach of avcsar2023comparative. This task is evaluated in both zero-shot and by fine-tuning the decoder, reporting Pearson correlation between predicted and true gene expression values. We explain these tasks in further details in Appendix D.2.
@@
Baselines. We benchmark HEIST against a diverse set of baselines, including graph-based spatial models, single-cell foundation models, and recent spatial foundation approaches. In particular, we include STAGATE (dong2022deciphering) and GraphST (long2023spatially), which capture local spatial relationships but struggle to generalize across datasets. We also evaluate against scFoundation (hao2023large), a large-scale single-cell foundation model that ignores spatial context. We also include comparisons with recent spatial foundation models CellPLM (wen2023cellplm), NoVAE (novae) and scGPT-spatial (wang2025scgpt), which incorporate spatial information but do not explicitly model hierarchical gene-cell interactions. Finally, for the gene imputation task, we compared against MAGIC a denoising method based on graph diffusion. For proteomics datasets such as UPMC, baselines like scGPT-spatial and CellPLM were originally pretrained on fixed gene vocabularies, supporting only 6 of the 28 markers available. To enable comparison, we report two settings: (i) unaligned, where models process only the subset of overlapping genes, and (ii) aligned, where we manually remap marker identities to the closest supported genes following prior work. It is important to note that scGPT pretrains on single-cell spatial trasncritomics data mixed with Visium (spot-level), while CellPLM combines scRNA-seq with single-cell spatial transcriptomics. HEIST, in contrast, is trained exclusively on single-cell resolution spatial transcriptomics. However, there is major overlap in the single-cell spatial transcriptomics portion of training data between HEIST and the baselines.
@@ 3.1 Results
Table 2: Performance on clinical outcome prediction. We classify cancer outcome, cancer remission, treatment response and placental conditions. Results on the right of vertical are proteomics. NEM stands for "not enough markers" error from Novae when it can not match enough markers. *Results from (wu2022space)
@@
Ablations. Table 4 shows that removing any key component of HEIST leads to performance degradation across tasks, confirming their importance. Hierarchical modeling and spatial information are most critical, with their removal causing the largest drops. Pre-training is crucial, especially for clinical outcome prediction due to the skewed label distribution in these datasets. Cross-level message passing and contrastive learning significantly improve cell classification, while the MAE is crucial for gene imputation and classification. In Table A5, we compare different positional encoding including Laplacian PEs, random walk based PEs, and sinusoidal PE. We notice that sinusoidal PEs outperform the other PE methods. In Table A6, we present the results with constant and adaptive $\tau$ (threshold for the co-expression graph), showing that the adaptive method we use outperforms the constant threshold. In Table A7, we ablate by removing each of the contrastive term; Results indicate that the three terms combinely capture complementary structural signals. We also present a way to interpret HEIST attention in Figure A4 and Table A8, indicating that HEIST attention focuses on niches and corresponds to known ligand-receptor pairs than existing models. In Table A9, we show that HEIST embeddings can predict LR pairs better than baselines. In Table A10, we show show different ways of mixing positional encodings and adding raw features. We notice that passing raw features through MLP before adding them results in performance drop in imputation because because the MLP-based mapping smooths or compresses the raw coordinate and expression signals in a way that removes fine-grained local variation. In Table A11, we show that using smaller batch sizes results in slight decrease in performance as smaller batch sizes provide less microenvironmental cues.
@@ Appendix A Related works
Hierarchical graph neural networks. Hierarchical graph neural networks operate on graphs with multiple levels, where each level represents a distinct granularity or type of information. MSMGN (fortunato2022multiscale), HOOD (Grigorev_2023_CVPR), and BSMS-GNN (pmlr-v202-cao23a) perform cross-level message passing across levels, but the levels are of same modality at different resolutions (e.g., coarsened versions of spatial graphs), rather than integrating graphs of distinct modalities and they are geared towards propagating signals globally. HIGH-PPI (gao2023hierarchical) introduces a hierarchical graph where one level encodes protein-protein interactions and the other encodes residue-level interactions. However, it does not include cross-level message passing, which limits the ability to combine information across biological levels. In contrast, HEIST works information from different modalities while integrating information across levels.
@@
Graph-based Spatial Transcriptomics methods. Graph-based spatial transcriptomics methods construct cell graphs based on spatial proximity and apply graph neural networks to model spatial relationships. SpaGCN (spagcn), CCST (ccst), STAGATE (dong2022deciphering), ConST (const), and GraphST (long2023spatially) train graph neural networks in un/self-supervised manner by using techniques such as deep graph infomax (dgi), iterative clustering (braun2022iterative), and auto-encodings task on gene expression (fang2024scmae). Unlike HEIST, these models do not generalize to other tissues without retraining and do not incorporate hierarchical modeling across genes and cells.
@@
scRNA-seq and Spatial Transcriptomics Foundation models. Foundation models for single-cell and spatial transcriptomics aim to pretrain general-purpose representations that can be transferred across datasets and tasks. scFoundation (hao2023large), geneformer (theodoris2023transfer), and scGPT (cui2024scgpt) extend transformer architectures to model scRNA-seq data by treating gene expression profiles as sequences. However, these models assume an ordering over genes and do not explicitly capture cell-cell relationships or spatial information. scGPT-spatial (wang2025scgpt) adapts scGPT to spatial transcriptomics. CellPLM (wen2023cellplm) incorporates cell-cell interactions in pertaining and adds a gaussian mixture prior to overcome data limitations. But CellPLM and scGPT-spatial operate over a fixed set of genes and do not model hierarchy. While transformer based models like scGPT-spatial can be considered as graph neural networks over fully conneceted graphs joshi2025transformers, they most biological networks are far from being fully connected, and explicitly modeling sparse, biologically plausible interactions is critical and enabled by graph-based learning. Existing foundation models thus either neglect spatial dependencies, gene co-expression structure, or both, limiting their ability to capture the context-specific nature of spatial transcriptomics. Furhtermore, their reliance in on gene embeddings hinders their generalization to proteomics data. In contrast, HEIST hierarchically models both this information and computes expressive and generalizable representations by working over gene co-expression networks.